OpenStudy (anonymous):

After reporting back to the park ranger about the food delivery schedule, you are then asked how much it costs to feed certain animals. You discover that the price to feed each lion is different than the price to feed each coyote. On Monday, the park was billed $315 to feed 21 coyotes and 21 lions. On Tuesday, the park was billed $484 to feed 28 coyotes and 36 lions. What is the cost to feed each lion and each coyote? You get a notification that next week, the prices to feed both animals will increase by 5%. What will be the cost to feed 35 coyotes and 40 lions?

OpenStudy (anonymous):

can someone please walk me through this I have no idea how to solve it.

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@DanJS @some.random.cool.kid

OpenStudy (solomonzelman):

it is a system of equations for the second paragraph.

OpenStudy (solomonzelman):

`On Monday, the park was billed $315 to feed 21 coyotes and 21 lions.` \(\large\color{black}{ 21C+216L=315 }\) `On Tuesday, the park was billed $484 to feed 28 coyotes and 36 lions. ` \(\large\color{black}{ 28C+36L=484 }\) where C is the cost per coyote, and L is a cost per lion.

OpenStudy (anonymous):

youre supposed to answer with the cost of feeding a lion and coyote from Monday and Tuesday and the cost of feeding the two animals combined the week after

OpenStudy (solomonzelman):

first you need to find the cost per each animal, no?

OpenStudy (anonymous):

yes, but im confused on how to do that. i know what the equation would look like but im unclear on how to solve it

OpenStudy (solomonzelman):

it is supposed to be \(\large\color{black}{ 21C+21L=315 \\ 28C+36L=484 }\) this is how it is.

OpenStudy (solomonzelman):

you can divide the first equation by 21, and multiply it times -28. After doing this, add the equations.

OpenStudy (anonymous):

-420? im sorry, I still don't get it.

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

\(\large\color{black}{ 21C+21L=315 \\ 28C+36L=484 }\) \(\large\color{black}{ C+L=15 \\ 28C+36L=484 }\) \(\large\color{black}{ -28C~~~-28L~~=~~-420 \\ 28C~~~~~~+36L~~=~~484 }\)

OpenStudy (solomonzelman):

can you add the equations? (here you can use my cde, if you want, ``` \(\large\color{black}{ -28C~~~-28L~~=~~-420 \\ 28C~~~~~~+36L~~=~~484 }\) ```

OpenStudy (solomonzelman):

just copy paste it, but you don't need, you can just type if you would like to.

OpenStudy (anonymous):

\[8L = 64\] \[L = 8\]